# CoAM: Corpus of All-Type Multiword Expressions

### v1

@@ A.1 Data Sources
Source Paper URL License Note
News Glavaš and Štajner (2013); Yimam et al. (2018) https://takelab.fer.hr/data/evsimplify/
https://sites.google.com/view/cwisharedtask2018/datasets
Creative Commons Attribution-NonCommercial-ShareAlike 3.0 We used the News_Train data from the CWI Shared Task 2018 datasets.
Commentary Kocmi et al. (2023) https://data.statmt.org/news-commentary/v18.1/training-monolingual/ Can be freely used for research purposes. We used the English part of v18.1.
TED (NAIST-NTT) Neubig et al. (2014) https://ahcweb01.naist.jp/old/resource/tedtreebank/ Creative Commons ShareAlike-Attribution-NonCommercial
TED (IWSLT) Cettolo et al. (2012, 2017) https://wit3.fbk.eu/2017-01 https://wit3.fbk.eu/2017-01-b Creative Commons Attribution-NonCommercial-NoDerivs 3.0
UD Bies et al. (2012); Walsh et al. (2018) https://gitlab.com/parseme/parseme_corpus_en Creative Commons ShareAlike 4.0 We used the portion of PARSEME corpus from the English Web Treebank.
Table 9: Data sources of CoAM.
A.2 MWE definition in Annotation Guidelines
  1. a.
    An MWE consists of at least two words that are always realized by the same lexemes. [^10] https://en.wikipedia.org/wiki/Lexeme For example, the MWE "break up" is always realized by (1) "break" or its conjugated form such as "broke" and (2) "up". Such words cannot be replaced without distorting the meaning of the expression or violating the language conventions.
  2. b.
    • Semantic idiomaticity occurs when the meaning of an expression cannot be derived from its components. That is, you cannot necessarily infer the meaning of the expression even if you know all the senses of its components. In other words, a semantically idiomatic takes on a meaning that is unique to that combination of words.
      The inferability differs from one expression to another. The meaning of idiomatic MWEs such as "kick the bucket" cannot be inferred from its components at all. The meaning of institutionalized phrases like "traffic light" is inferable to some degree. Yet it is an MWE because it does not mean any type of light related to traffic but a specific type of light.
  3. c.
    An MWE is not a proper noun, such as the name of a specific person, organization, and so forth. In this project, do not annotate proper nouns unless they are part of an MWE (like Pavlov in "Pavlov’s dog"). Here is the full list of proper nouns in our definition.111111 The 11 types derive from the named entity types of OntoNotes Release 5.0 Weischedel et al. (2013).
    Type Example
    People’s names Shohei Ohtani
    Facilities Narita International Airport, Taipei 101
    Organizations Procter & Gamble, Kyoto University
    Geopolitical entities (GPE) Los Angeles, the United Kingdom
    Non-GPE locations Mount Fuji, Amazon River
    Products Toyota Prius, Samsung Galaxy
    Named events World War II, Olympic Games Tokyo 2020
    Works of art Norwegian Wood, Bohemian Rhapsody
    Named legal documents The Magna Carta
    Named languages Middle English, American Sign Language
    Note:
    • Proper noun phrases usually start with a capital letter, while exceptions like "iPhone 15" exist.
A.3 Notes on Idiomaticity and (Non-)Compositionality
@@
To judge whether a sequence is an MWE, we focus on (semantic) idiomaticity instead of non-compositionality, although non-compositionality has been considered an inherent property of MWEs in previous studies, such as Tedeschi et al. (2022). To discuss the difference between the two notions, consider the expression spill the beans. It is deemed compositional because it can be analyzed as being made up of spill in a “reveal” sense and the beans in a “secret(s)” sense, resulting in the overall compositional reading of “reveal the secret(s)” Sag et al. (2002). Meanwhile, the “secret(s)” sense is unique to the expression and cannot be derived from the word beans, making the whole expression semantically idiomatic. We argue that spill the beans should be identified as an MWE because the meaning of beans depends on the whole expression and that any idiomatic sequences should be identified even if they are compositional.
A.4 Annotation Interface
(a) An example of discontinuous MWE.
(b) An example of overlapping MWEs.
Figure 2: How to annotate tricky MWEs using our interface.
Our interface generated by CAIGen enables annotators to mark spans in any form, including discontinuous and overlapping ones, only by checking checkboxes. Figure 2 shows the examples.
@@ B.1 LLM fine-tuning
Learning rate 2e-4
@@
Long Llama-8B 29.4±2.1 $82.6$±1.2 17.9±1.6 5.9±0.8 26.4±1.5 24.2±3.5 0.0±0.0
Qwen-7B 45.2±0.6 63.2±0.9 35.2±0.8 26.0±2.0 47.2±1.0 31.8±2.0 0.0±0.0
Short Llama-8B 8.6±14.8 28.9±50.0 5.0±8.7 1.4±2.4 7.1±12.4 7.2±12.5 $\mathbf{0.0}$±0.0
Qwen-7B 32.8±0.2 $70.8$±1.6 21.4±0.1 12.1±0.5 30.5±0.6 22.3±0.7 0.0±0.0
@@
To validate the efficacy of the long MWE definition, we perform an ablation study comparing the long definition to the short definition. They are both a summary of the full definition described in Section A.2. As shown in Table 12, the long definition has 162 words while the short is further contracted to 57 words. We perform experiments with Llama-8B and Qwen-7B, using the tsv_to_tsv format and the hyperparameters described in Table 10 (same as the main experiments).
@@ Appendix C Computational Budgets
In experiments for MaW (Rule+BiEnc and Rule+DCA), we use a single NVIDIA RTX 2080Ti GPU with 11GB RAM. Each run of training and testing takes around 8 minutes. Consequently, the total GPU hours for MaW are estimated to be 0.8 hours.
In LLM fine-tuning, we use NVIDIA RTX 6000 GPU with 48GB RAM for smaller models (Llama-8B and Qwen-7B) and NVIDIA A100 PCIe with 80GB RAM for larger models (Llama-70B and Qwen-72B). Each run of training and testing takes around 66 minutes for smaller models and 588 minutes for larger models. Thus, the total GPU hours for fine-tuning experiments of smaller and larger models are estimated to be 6.6 hours and 58.8 hours, respectively.

### v2

@@ A.1 Data Sources
See Table 9.
A.2 Notes on Idiomaticity and (Non-)Compositionality
@@
To judge whether a sequence is an MWE, we focus on (semantic) idiomaticity instead of non-compositionality, although non-compositionality has been considered an inherent property of MWEs in previous studies, such as Tedeschi et al. (2022). To discuss the difference between the two notions, let us consider the expression spill the beans. It is deemed compositional because it can be analyzed as being made up of spill in a “reveal” sense and the beans in a “secret(s)” sense, resulting in the overall compositional reading of “reveal the secret(s)” Sag et al. (2002). Meanwhile, the “secret(s)” sense is unique to the expression and cannot be derived from the word beans, making the whole expression semantically idiomatic. We argue that spill the beans should be identified as an MWE because the meaning of beans depends on the whole expression, and that any idiomatic sequences of this kind should be identified even if they are compositional.
Algorithm 1 Algorithm of MWE Type Tagging
1:MWE $\bm{m}$
2:Type $t$
3:for all $w_{i}\in\bm{m}$ do
4:     if $w_{i}=w^{*}$ then
5:         if PoS of $w_{i}\in\{\text{NOUN},\text{PRON},\text{PROPN}\}$ then
6:              if children of $w_{i}$ include a relative clause modifier that is verb then
7:                  $t\leftarrow$ Verb $\triangleright$ E.g., $\ldots$ the price he pays $\ldots$
8:              end if
9:              $t\leftarrow$ Noun $\triangleright$ E.g., $\ldots$ born in the middle of nowhere $\ldots$
10:         else if PoS of $w_{i}\in\{\text{VERB},\text{AUX}\}$ then
11:              if nominal subject of $w_{i}\in\bm{m}$ then
12:                  $t\leftarrow$ Clause $\triangleright$ E.g., $\ldots$ when it comes to climate change $\ldots$
13:              else
14:                  $t\leftarrow$ Verb $\triangleright$ E.g., $\ldots$ the expression stands for $\ldots$
15:              end if
16:         else if PoS of $w_{i}\in\{\text{ADP},\text{ADJ},\text{ADV},\text{CCONJ},\text{SCONJ}\}$ then
17:              $t\leftarrow$ Mod/Conn $\triangleright$ E.g., $\ldots$ for at least two decades $\ldots$
18:         else
19:              $t\leftarrow$ Other (other PoS) $\triangleright$ E.g., $\ldots$ de facto target gauge $\ldots$
20:         end if
21:     else
22:         $t\leftarrow$ Other (head not in MWE) $\triangleright$ E.g., $\ldots$ , and so on $\ldots$
23:     end if
24:end for
25:return $t$
A.3 Algorithm for MWE Type Tagging
We tag MWEs in CoAM by the algorithm shown in Algorithm 1.
MWE Type Discontinuous (%) Example
Noun 02.0 …a suicide car bomber and Taliban militants …
Verb 21.8 turned four aircraft into cruise missiles …
Mod/Conn 04.8 …concentration of power in his own hands.
Clause 00.0
Other 20.8 no more than 2.7%.
Table 10: Ratio of discontinuous MWEs in CoAM by MWE type.
A.4 Statistics Regarding MWE Continuity
Discontinuous MWEs are common in English, and previous work has stressed the importance of identifying them (Schneider et al., 2014a; Rohanian et al., 2019). Table 10 shows the ratio of discontinuous MWEs by MWE type. Verb MWEs have the highest proportion of discontinuous expressions, while Mod/Conn, Noun, and Other MWEs reach lower values.
@@ B.1 LLM fine-tuning
Learning rate 1e-4 for Llama-70B 2e-4 for other models
@@
Long Llama-8B 24.9±3.7 92.0±2.0 14.4±2.4 3.3±0.0 19.4±2.9 21.6±4.8 0.0±0.0
Qwen-7B 48.1±1.0 60.9±0.8 39.7±1.0 28.7±1.3 52.0±0.8 39.6±0.9 0.0±0.0
Short Llama-8B 21.0±3.1 89.9±4.7 11.9±2.0 3.3±0.0 16.3±2.3 16.8±5.0 0.0±0.0
Qwen-7B 36.7±1.3 65.7±0.7 25.5±1.3 15.4±1.7 32.6±2.2 29.7±0.0 0.0±0.0
@@
To validate the efficacy of the long MWE definition, we perform an ablation study comparing the long definition to the short definition. They are both a summary of the full definition described in Appendix D. As shown in Table 13, the long definition has 162 words while the short is further contracted to 57 words. We perform experiments with Llama-8B and Qwen-7B, using the tsv_to_tsv format and the hyperparameters described in Table 11 (same as the main experiments).
@@ Appendix C Computational Budgets
For experiments for MaW (Rule+BiEnc and Rule+DCA), we use a single NVIDIA RTX 2080Ti GPU with 11GB RAM. Each run of training and testing takes around 8 minutes. Consequently, the total GPU hours for MaW are estimated to be 0.8 hours.
For LLM fine-tuning, we use NVIDIA RTX 6000 GPU with 48GB RAM for smaller models (Llama-8B and Qwen-7B) and NVIDIA A100 PCIe with 80GB RAM for larger models (Llama-70B and Qwen-72B). Each run of training and testing takes around 66 minutes for the smaller models and 588 minutes for the larger models. Thus, the total GPU hours for fine-tuning experiments of smaller and larger models are estimated to be 6.6 hours and 58.8 hours, respectively.
Appendix D Excerpt of Annotation Guidelines
D.1 Definition of MWEs
  1. a.
    An MWE consists of at least two words that are always realized by the same lexemes. [^12] Lexeme is a set of words related through inflection. Words belonging to the same lexeme share a common lemma. For example, the MWE "break up" is always realized by (1) "break" or its conjugated form such as "broke" and (2) "up". Such words cannot be replaced without distorting the meaning of the expression or violating the language conventions.
  2. b.
    • Semantic idiomaticity occurs when the meaning of an expression cannot be derived from its components. That is, you cannot necessarily infer the meaning of the expression even if you know all the senses of its components. In other words, a semantically idiomatic expression takes on a meaning that is unique to that combination of words.
      The inferability differs from one expression to another. The meaning of idiomatic MWEs such as "kick the bucket" cannot be inferred from its components at all. The meaning of institutionalized phrases like "traffic light" is inferable to some degree. Yet "traffic light" is semantically idiomatic because it does not mean any type of light related to traffic but a specific type of light.
  3. c.
    An MWE is not a proper noun, such as the name of a specific person, organization, and so forth. In this project, do not annotate proper nouns unless they are part of an MWE (like Pavlov in "Pavlov’s dog"). Proper nouns usually start with a capital letter, while exceptions like "iPhone 15" exist. Below is the full list of proper nouns in our definition.131313 The 11 types derive from the named entity types of OntoNotes Release 5.0 Weischedel et al. (2013).
    Type Example
    People’s names Shohei Ohtani
    Facilities Narita International Airport, Taipei 101
    Organizations Procter & Gamble, Kyoto University
    Geopolitical entities (GPE) Los Angeles, the United Kingdom
    Non-GPE locations Mount Fuji, Amazon River
    Products Toyota Prius, Samsung Galaxy
    Named events World War II, Olympic Games Tokyo 2020
    Works of art Norwegian Wood, Bohemian Rhapsody
    Named legal documents The Magna Carta
    Named languages Middle English, American Sign Language
D.2 Notes
MWEs containing replaceable words
MWEs are not necessarily made of continuous words. Some MWEs contain open slots, that is, words that may be replaced with a large or open class of words. In annotation, we do not include open slots in MWEs, as illustrated by the following examples:
  • You took me by surprise!
  • Pick me up at the station.
Note that if a word is replaceable with only a very small number of alternatives, we consider it as part of the MWE. For example, we count the following as different MWEs:
  • Their food leaves a lot to be desired.
  • The work leaves much to be desired.
Overlapping MWEs
There is a chance that multiple MWEs share the same word. For example, letting in the following sentence belongs to both letting in and letting out.
Rearranged MWEs
There is a chance that the order of component words in an MWE is different from its canonical form, but we still count such rearranged sequences as MWEs. In the following example, rearrangement happens in the MWE break <one’s> heart. As a side note, my here is not annotated because it is replaceable.
Hyphens
On our annotation interface, hyphens are treated as words. When you encounter a hyphen in an MWE, mark the hyphen. Meanwhile, note that some hyphenated words, e.g., twenty-one are not MWEs and should not be annotated.